OpenStudy (anonymous):

What is the solution to y = 2x + 3 for the replacement set x = {1.3}?

OpenStudy (palwandew):

Simple. Substitue "1.3" for x in the given equation y = 2x + 3. 2x + 3 = y 2 * 1.3 + 3 = y 2.6 + 3 = y 5.6 = y The solution to the equation with the replacement set {1.3} for x is {5.6} = y. I hope you understand how to solve the equation! :)

OpenStudy (anonymous):

so the answer is .

OpenStudy (palwandew):

Y= {5.6}

OpenStudy (anonymous):

ok thax